OpenStudy (anonymous):

A new 13.3'' Macbook Pro costs $1,200.00, and a one year old 13.3'' Macbook Pro costs $1,000.00. How do I make an exponential decay model with this?

OpenStudy (australopithecus):

Do you have the exponential decay formula I have forgotten it

OpenStudy (australopithecus):

I jsut need to see the formula I dont need to know what the variables mean

OpenStudy (anonymous):

I think it's y=a(1-r)t

OpenStudy (australopithecus):

Ugh deleted what I wrote. D = Pe^(rt) P =1200.00 D=1000.00 t = 1 year r = ? (Rate) solve for r

OpenStudy (australopithecus):

I assume you know how to use logirthms to solve this

OpenStudy (australopithecus):

oops the equation is D = Pe^(-rt)

OpenStudy (anonymous):

I forgot what that is.

OpenStudy (australopithecus):

ln(D/P) = ln(e^(-rt))

OpenStudy (australopithecus):

does that help?

OpenStudy (anonymous):

sorry but no it doesn't.

OpenStudy (australopithecus):

Logrithmic rules 1) log_b(mn) = log_b(m) + logb(n) 2) log_b(m/n) = log_b(m) – logb(n) 3) log_b(m^(n)) = n · log_b(m) 4) log_3(3) = 1 5) log_e(x) = ln(x) 6) log(x) = y is the same as x =10^(y) 7) log_10(x) = log(x)

OpenStudy (australopithecus):

Does that help?

OpenStudy (australopithecus):

Log_3(x) =\[\log_3(x)\]

OpenStudy (anonymous):

Yes it does. Thanks.

OpenStudy (australopithecus):

if you want me to check your answer I can do that as well

OpenStudy (anonymous):

no it's okay.

OpenStudy (australopithecus):

ok then back to the grind for me